y=3x+2
6x-2y=-4
pick between (3,0),(-3,1),infinitely many solutions, or no solution

Answer by mananth(16946)

y=3x+2
.....
6x-2y=-4
/2
3x-y=-2
y=3x+2
..
the equations are of the same line.
so there are infintely many solutions
...
m.ananth@hotmail.ca